OpenStudy (anonymous):

Solve: Sq Root 3x=4-3=7

OpenStudy (anonymous):

\[\sqrt{3x+4}-3=7\]

OpenStudy (anonymous):

\[\sqrt{3x+4}=11------------>3x+4=121\] rest it's easy

OpenStudy (anonymous):

why 11?

OpenStudy (anonymous):

sry 10. :)

OpenStudy (anonymous):

and the check should equal 7?

OpenStudy (anonymous):

3x=96 x=32

OpenStudy (anonymous):

thats what I got but the check wont go back to 7

OpenStudy (anonymous):

is it sqrt(3x+4)-3 or sqrt(3x+4-3) ?

OpenStudy (anonymous):

just the first one

OpenStudy (anonymous):

\[\sqrt{3x+4}-3=7\]

OpenStudy (anonymous):

3x+4=100 3x=96 x=32 don't know what esle to say....

OpenStudy (anonymous):

ty